1.      A company is offering a job with a salary of Rs. 50,000.00 for the first year and a 4% raise each year after that. If that 4% raise continues every year,

a)        Find the total amount of money an employee would earn in a 10-years career.


1
Expert's answer
2021-06-17T05:00:29-0400

The first year: "A_1=50000."

The growth factor: "r=0.04."

We have the geometric progression


"A_1, A_1(1+r), A_1(1+r)^2, ..."

"S_{10}=\dfrac{A_1(1-(1+r)^{10})}{1-(1+r)}"

"=\dfrac{A_1((1+r)^{10})-1}{r}"

"=\dfrac{50000((1+0.04)^{10}-1)}{0.04}=Rs\ 600305.36"

An employee would earn Rs. 600305.36 in a 10-years career.